# Online Fair Allocations with Binary Valuations and Beyond

### v1

@@ 3 Allocation of Goods
In this section, we focus on the setting of goods. Here are some definitions of fairness and efficiency notions.
@@
For any constant $\alpha\in[0,1]$, an allocation $A^{k}$ is $\alpha$-approximate envy-free up to one good ($\alpha$-EF1) if, for every pair of agents $i,i^{\prime}\in N$ with $A_{i^{\prime}}^{k}\neq\emptyset$, it holds that $v_{i}(A_{i}^{k})\geq\alpha\cdot v_{i}(A_{i^{\prime}}^{k}\setminus\{e\})$ for some $e\in A_{i^{\prime}}^{k}$.
@@
where $\Pi_{n}(T^{k})$ is the set of allocations of $T^{k}$. For any constant $\alpha\in[0,1]$, an allocation $A^{k}$ is $\alpha$-approximate maximin share fair ($\alpha$-MMS) if for any agent $i\in N$, it holds that $v_{i}(A_{i}^{k})\geq\alpha\cdot\mathsf{MMS}_{i}^{k}$.
@@
The utilitarian social welfare $USW$ of allocation $A^{k}$ is given by $USW(A^{k})=\sum_{i\in N}v_{i}(A_{i}^{k})$. For any constant $\alpha\in[0,1]$, an allocation $A^{k}$ is $\alpha$-max-USW if $USW(A^{k})\geq\alpha\cdot\max_{B^{k}\in\Pi_{n}(T^{k})}USW(B^{k})$ holds, where $\Pi_{n}(T^{k})$ is the set of allocations of $T^{k}$.
In addition to the above definitions, we also pay attention to the following restriction, which is added in the goods setting.
@@
An allocation $A^{k}$ is non-wasteful (NW) if (1) $\Delta_{A_{i}^{k}\setminus\{e\}}(e)>0$ for any $i\in N$ and any good $e\in A^{k}_{i}$. (2) $\Delta_{A_{i}^{k}}({e}^{{}^{\prime}})=0$ for any discarded good $e^{\prime}\in T^{k}$.
When agents have additive valuation functions, for any arriving good $e$, as long as it is allocated to one agent who values it positively, an NW allocation can be found. Next, We explore the existence of submodular valuation functions.
Claim 1.
@@
Given an instance $(T,N,(v_{i})_{i\in N})$, Algorithm 1 can compute an NW allocation in polynomial time. When item $e_{k}$ ($k\in[t]$) arrives, if there exists an agent $j\in N$ such that $\Delta_{e_{k}}^{j}(A_{j}^{k-1})=1$, $e_{k}$ will be allocated. We assume that $e_{k}$ is allocated to agent $i$. We will show that $\Delta_{e_{k}}^{i}(A_{i}^{r}\setminus\{e_{r}\})=1$ holds for any round $r>k$ after some new items are added to agent $i$’s bundle. Without loss of generality, we assume that $e_{s}$ is an arbitrarily newly added item in agent $i$’s bundle, where $s>k$, it holds that $\Delta_{e_{s}}^{i}(A_{i}^{s-1})=v_{i}(A_{i}^{s-1}\cup\{e_{s}\})-v_{i}(A_{i}^{s-1})=1$. Then, we have
$\Delta_{e_{k}}^{i}(A_{i}^{s}\setminus\{e_{r}\})$ $=v_{i}(A_{i}^{s-1}\cup\{e_{s}\})-v_{i}(A_{i}^{s-1}\cup\{e_{s}\}\setminus\{e_{k}\})$
$=v_{i}(A_{i}^{s-1})+1-v_{i}(A_{i}^{s-1}\cup\{e_{s}\}\setminus\{e_{k}\})$
$\geq v_{i}(A_{i}^{s-1})+1-v_{i}(A_{i}^{s-1}\setminus\{e_{k}\})-v_{i}(e_{s})$
$=v_{i}(A_{i}^{s-1})-v_{i}(A^{s-1}\setminus\{e_{k}\})$
$=\Delta_{e_{k}}^{i}(A_{i}^{s-1}\setminus\{e_{k}\}),$
where the inequality holds for submodularity. Following this induction by repeatedly utilizing submodularity, we have that $\Delta_{e_{k}}^{i}(A_{i}^{s}\setminus\{e_{k}\})\geq\cdots\geq\Delta_{e_{k}}^{i}(A_{i}^{k}\setminus\{e_{k}\})=\Delta_{e_{k}}^{i}(A_{i}^{k-1})=1$. Since $v_{i}$ is a submodular binary valuation function, it holds that $\Delta_{e_{k}}^{i}(A_{i}^{s}\setminus\{e_{k}\})\leq 1$. Thus, we can conclude that $\Delta_{e_{k}}^{i}(A_{i}^{s}\setminus\{e_{k}\})=1$. ∎
Claim 2.
Proof.
We construct an instance with two agents, $v_{1}=v_{2}=v$. When first item $e_{1}$ arrives, $v(e_{1})=3$. W.l.o.g., we allocate $e_{1}$ to agent 1. When the second item $e_{2}$ arrives, $v(e_{2})=3,v(e_{1},e_{2})=5$. Case (1): we allocate $e_{2}$ to agent 2, for the third item $e_{3}$,$v(e_{3})=5,v(e_{1},e_{3})=v(e_{2},e_{3})=v(e_{1},e_{2},e_{3})=5$. We have that $v(e_{1},e_{3})-v(e_{3})=v(e_{2},e_{3})-v(e_{3})=0$. So, no matter who is assigned $e_{3}$, it cannot satisfy NW. Case (2): we allocate $e_{2}$ to agent 1, for the third item $e_{3}$ with $v(e_{3})=5,v(e_{1},e_{3})=v(e_{2},e_{3})=v(e_{1},e_{2},e_{3})=5$. To ensure NW, we must allocate $e_{3}$ to agent 2. For the fourth item $e_{4}$, $v(e_{4})=v(e_{1},e_{4})=v(e_{2},e_{4})=v(e_{3},e_{4})=v(e_{1},e_{2},e_{4})=v(e_{1},e_{3},e_{4})=v(e_{2},e_{3},e_{4})=v(e_{1},e_{2},e_{3},e_{4})=6$. We have that $v(e_{1},e_{2},e_{4})-v(e_{2},e_{4})=v(e_{3},e_{4})-v(e_{4})=0$. So, no matter who is assigned $e_{4}$, it cannot satisfy NW. ∎
@@
Output: A non-wasteful allocation $A$
1 Initialize: Let $A=(\emptyset,\ldots,\emptyset)$;
@@ 3.1 The Impossibilities of Approximate EF1 and MMS
In this part, we present a strong negative result regarding the approximations of EF1 and MMS for general additive valuations. The valuation function $v_{i}$ is tri-valued if $v_{i}(e)\in\{a_{i},b_{i},z_{i}\}$ for any $e\in T$, where $0<a_{i}\leq b_{i}\leq z_{i}$.
@@
For general additive functions, no deterministic online algorithm can compute an NW allocation that guarantees $\alpha$-EF1 or $\alpha$-MMS for any constant $\alpha>0$, even for two agents with additive tri-valued valuation functions.
@@
Consider an online instance with two agents. For any agent $i\in N$ and item $e\in T$, $v_{i}(e)\in\{\epsilon,1,\frac{1}{\epsilon}\}$, where $10<\epsilon<$ 1. The items arrive online, and the valuation of each item is shown in Table 3. When $t=1$, the arriving item has the value of 1 for both agents. Without loss of generality, agent 1 picks it. When $t=2$, the item that has the value of $\frac{1}{\epsilon}$ for agent 1 and $\epsilon$ for agent 2 arrives. In order to avoid 00-EF1 or 00-MMS, this item must be allocated to agent 2. In round $t=3$, the arriving item has the value of $\frac{1}{\epsilon}$ for both agents. In this case, if this item is allocated to agent 1, for agent 2, the approximation ratio of EF1 is $\epsilon$, and the approximation ratio of MMS is $\frac{\epsilon}{1+\epsilon}$; if this item is allocated to agent 2, for agent 1, the approximation ratio of EF1 is $\frac{1}{\frac{1}{\epsilon}}=\epsilon$ and the approximation ratio of MMS is $\frac{1}{1+\frac{1}{\epsilon}}=\frac{\epsilon}{1+\epsilon}$. Therefore, for any constant $\alpha>0$, let $\epsilon<\alpha$, and we can conclude that no algorithm can guarantee an NW allocation that is $\alpha$-EF1 or $\alpha$-MMS for any given $\alpha>0$. This result can be generalized to more than two agents, and we defer the proof to Proposition 5.
$t=1$ $t=2$ $t=3$
@@
Agent 2 $1$ $\epsilon$ $\frac{1}{\epsilon}$
Table 3: The impossibility results of EF1 and MMS for additive tri-valued valuations
By Theorem 1, we cannot make further progress in the approximation of EF1 and MMS for tri-valued valuation functions and beyond in the online allocation problem. This means that we can only focus our efforts on bi-values. In the following parts, we focus on two classes of valuation functions, including submodular binary and additive bi-valued valuation functions, and study whether there are positive results about the approximation of EF1 or MMS.
@@ 3.2 Submodular Binary Valuations
This part focuses on submodular binary valuation functions, where for each agent $i\in N$, her valuation function $v_{i}$ is submodular and the marginal value of any item is 0 or 1. For a subclass of submodular binary valuation functions based on maximum (unweighted) bipartite matching, which is also called assignment [19] or OXS [20] valuations, Hosseini et al.[11] show that online allocations satisfying $\frac{1}{2}$-EF1 and $\frac{1}{2}$-MMS exist. In their model, there are some groups, each of which includes some members, and each member can only receive one item at most. When an item arrives, they greedily match the item to some member who likes the item and has never received any items. Leaving the online matching model, we consider the general online allocation problem, greedily assigning each arriving item to the agent that brings the maximum marginal utility. Deriving from the greedy technique of the submodular optimization problem and the sequential allocating technique to guarantee fairness in [11], we propose an algorithm named Marginal-Greedy Algorithm (Algorithm 2) for general binary valuation functions. We prove that input submodular binary valuations into Algorithm 2, we can get an NW allocation that is $\frac{1}{2}$-EF1, $\frac{1}{2}$-MMS, and $\frac{1}{2}$-max-USW. Additionally, we conjecture that the Algorithm 2 will also perform well when we input subadditive binary valuation functions into Algorithm 2, and leave this conjecture as an open question.
Input: An instance $(T,N,(v_{i})_{i\in N})$ with general binay valuation functions
Output: A non-wasteful approximate EF1 and approximate MMS allocation $A$
1 Initialize: let $\pi=(\pi_{1},\cdots,\pi_{n})$ be an arbitrary order of $n$ agents;
2Let $A=(\emptyset,\ldots,\emptyset)$;

### v2

@@ 3 Allocation of Goods
Throughout this section, we focus on the goods setting. We first introduce some definitions of fairness and efficiency notions that are used in this section.
@@
For any $\alpha\in[0,1]$, an allocation ${\bf A}^{k}$ is $\alpha$-approximate envy-free up to one good ($\alpha$-EF1) if, for every pair of agents $i,i^{\prime}\in N$ with $A_{i^{\prime}}^{k}\neq\emptyset$, it holds that $v_{i}(A_{i}^{k})\geq\alpha\cdot v_{i}(A_{i^{\prime}}^{k}\setminus\{e\})$ for some $e\in A_{i^{\prime}}^{k}$.
@@
For any $\alpha\in[0,1]$, an allocation ${\bf A}^{k}$ is $\alpha$-approximate maximin share fair ($\alpha$-MMS) if for any agent $i\in N$, it holds that $v_{i}(A_{i}^{k})\geq\alpha\cdot\mathsf{MMS}_{i}^{k}$.
@@
The utilitarian social welfare of allocation ${\bf A}^{k}$ is given by $USW({\bf A}^{k})=\sum_{i\in N}v_{i}(A_{i}^{k})$. For any $\alpha\in[0,1]$, an allocation ${\bf A}^{k}$ is $\alpha$-max-USW if $USW({\bf A}^{k})\geq\alpha\cdot\max_{{\bf B}^{k}\in\Pi_{n}(T^{k})}USW({\bf B}^{k})$.
Besides the above definitions, we also pay attention to the following restriction, which is added in the goods setting.
@@
An allocation ${\bf A}^{k}$ is non-wasteful (NW) if (1) $\Delta_{A_{i}^{k}\setminus\{e\}}^{i}(e)>0$ for any agent $i\in N$ and any good $e\in A^{k}_{i}$, and (2) $\Delta_{A_{i}^{k}}^{i}({e}^{{}^{\prime}})=0$ for any agent $i\in N$ and any discarded good $e^{\prime}\in T^{k}$.
Note that imposing NW in the goods setting is necessary, as our online model permits item discarding (see our example for illustration). Without this constraint, achieving EF1 would be trivial by simply discarding all online items. For agents with additive valuation functions, an NW allocation can always be found as long as each arriving item is allocated to at least one agent who values it positively. Below, we present a positive result where an NW allocation always exists for submodular binary valuation functions that we study in the subsequent parts, and complement it with a negative result for general submodular valuation functions.
Lemma 1.
@@
Output: A NW allocation $\bf A$}
1 Initialize: Let ${\bf A}=(\emptyset,\ldots,\emptyset)$;
@@ 8        Let $A_{i}=A_{i}\cup\{e\}$;
Given an instance $(T,N,(v_{i})_{i\in N})$, Algorithm 1 can compute an NW allocation. When item $e_{k}$ ($k\in[t]$) arrives, if there exists an agent $j\in N$ such that $\Delta_{A_{j}^{k-1}}^{j}({e_{k}})=1$, $e_{k}$ will be allocated. We assume that $e_{k}$ is allocated to agent $i$. We will show that $\Delta_{A_{i}^{r}\setminus\{e_{r}\}}^{i}({e_{k}})=1$ holds for any round $r>k$ after some new items are added to agent $i$’s bundle. Without loss of generality, we assume that $e_{s}$ is an arbitrarily newly added item in agent $i$’s bundle, where $s>k$, it holds that $\Delta_{A_{i}^{s-1}}^{i}({e_{s}})=v_{i}(A_{i}^{s-1}\cup\{e_{s}\})-v_{i}(A_{i}^{s-1})=1$. Then, we have
$\Delta_{{A_{i}^{s}}\setminus\{e_{r}\}}^{i}({e_{k}})$ $=v_{i}(A_{i}^{s-1}\cup\{e_{s}\})-v_{i}(A_{i}^{s-1}\cup\{e_{s}\}\setminus\{e_{k}\})$
$=v_{i}(A_{i}^{s-1})+1-v_{i}(A_{i}^{s-1}\cup\{e_{s}\}\setminus\{e_{k}\})$
$\geq v_{i}(A_{i}^{s-1})+1-v_{i}(A_{i}^{s-1}\setminus\{e_{k}\})-v_{i}(e_{s})$
$=v_{i}(A_{i}^{s-1})-v_{i}(A^{s-1}\setminus\{e_{k}\})$
$=\Delta_{A_{i}^{s-1}\setminus\{e_{k}\}}^{i}({e_{k}}),$
where the inequality holds for submodularity. Following this induction by repeatedly utilizing submodularity, we have that $\Delta_{A_{i}^{s}\setminus\{e_{k}\}}^{i}({e_{k}})\geq\ldots\geq\Delta_{A_{i}^{k}\setminus\{e_{k}\}}^{i}({e_{k}})=\Delta_{A_{i}^{k-1}}^{i}({e_{k}})=1$. Since $v_{i}$ is a submodular binary valuation function, it holds that $\Delta_{A_{i}^{s}\setminus\{e_{k}\}}^{i}({e_{k}})\leq 1$. Thus, we can conclude that $\Delta_{A_{i}^{s}\setminus\{e_{k}\}}^{i}({e_{k}})=1$. ∎
Lemma 2.
Proof.
We construct an instance with two agents, $v_{1}=v_{2}=v$. When first item $e_{1}$ arrives, $v(e_{1})=3$. W.l.o.g., we allocate $e_{1}$ to agent 1. When the second item $e_{2}$ arrives, $v(e_{2})=3,v(\{e_{1},e_{2}\})=5$. Case (1): we allocate $e_{2}$ to agent 2, for the third item $e_{3}$$v(e_{3})=5,v(\{e_{1},e_{3}\})=v(\{e_{2},e_{3}\})=v(\{e_{1},e_{2},e_{3}\})=5$. We have that $v(\{e_{1},e_{3}\})-v(e_{3})=v(\{e_{2},e_{3}\})-v(e_{3})=0$. So, no matter who is assigned $e_{3}$, it cannot satisfy NW. Case (2): we allocate $e_{2}$ to agent 1, for the third item $e_{3}$ with $v(e_{3})=5,v(\{e_{1},e_{3}\})=v(\{e_{2},e_{3}\})=v(\{e_{1},e_{2},e_{3}\})=5$. To ensure NW, we must allocate $e_{3}$ to agent 2. For the fourth item $e_{4}$, $v(e_{4})=v(\{e_{1},e_{4}\})=v(\{e_{2},e_{4}\})=v(\{e_{3},e_{4}\})=v(\{e_{1},e_{2},e_{4}\})=v(\{e_{1},e_{3},e_{4}\})=v(\{e_{2},e_{3},e_{4}\})=v(\{e_{1},e_{2},e_{3},e_{4}\})=6$. We have that $v(\{e_{1},e_{2},e_{4}\})-v(\{e_{2},e_{4}\})=v(\{e_{3},e_{4}\})-v(e_{4})=0$. So, no matter who is assigned $e_{4}$, it cannot be NW. ∎
@@ 3.1 The Impossibilities of Approximate EF1 and MMS
(Aleksandrov et al., 2015) show that in the online setting, an EF1 allocation exists for additive binary valuations, where for each agent $i\in N$, $v_{i}$ is additive and $v_{i}(e)\in\{0,1\}$ for any $e\in T$. In this part, we present a strong negative result regarding the approximations of EF1 and MMS for general additive valuations. We say that the valuation function $v_{i}$ is additive personalized tri-valued if $v_{i}$ is additive and $v_{i}(e)\in\{a_{i},b_{i},z_{i}\}$ for any $e\in T$, where $0<a_{i}\leq b_{i}\leq z_{i}$.
@@
No deterministic online algorithm can compute an NW allocation that guarantees $\alpha$-EF1 or $\alpha$-MMS for any $\alpha>0$, even for two agents with additive personalized tri-valued valuation functions.
@@
$e_{1}$ $e_{2}$ $e_{3}$
@@
agent 2 11 $\epsilon$ $\frac{1}{\epsilon}$
Table 4: The impossibility result of EF1 and MMS.
Given any $\alpha>0$, let $\epsilon$ be a constant arbitrarily close to 0 such that $0<\epsilon<\alpha$. Consider an online instance with two agents. For any agent $i\in N$ and item $e\in T$, $v_{i}(e)\in\{\epsilon,1,\frac{1}{\epsilon}\}$, where $0<\epsilon<1$. The items arrive online, and the valuation of each item is shown in Table 4. When $t=1$, the item $e_{1}$ arrives, $T^{1}=\{e_{1}\}$. Without loss of generality, suppose it is allocated to agent 1, $A^{1}_{1}=\{e_{1}\},A^{1}_{2}=\emptyset$. When $t=2$, the item $e_{2}$ arrives, $T^{2}=\{e_{1},e_{2}\}$. To maintain the fairness of allocation, we must allocate it to agent 2, i.e., $A^{2}_{1}=\{e_{1}\},A^{2}_{2}=\{e_{2}\}$. Otherwise, agent 1 gets $\{e_{1},e_{2}\}$ and agent 2 gets an empty set, $v_{2}(\emptyset)=0<v_{2}(\{e_{1},e_{2}\}\setminus\{e\})$ for any $e\in\{e_{1},e_{2}\}$ and $MMS_{2}(T^{2})=\epsilon$. It is 0-EF1 and 0-MMS. When $t=3$, item $e_{3}$ arrives, $T^{3}=\{e_{1},e_{2},e_{3}\}$. $MS_{1}(T^{3})=\frac{1}{\epsilon}, MM$$MS_{2}(T^{3})=1+\epsilon$. If we allocate it to agent 1, i.e., $A^{3}_{1}=\{e_{1},e_{3}\},A^{3}_{2}=\{e_{2}\}$, we have $v_{2}(A^{3}_{2})=\epsilon$, $v_{2}(A^{3}_{1}\setminus\{e_{1}\})=\frac{1}{\epsilon}$, and $v_{2}(A^{3}_{1}\setminus\{e_{3}\})=1$. Thus, $\bf A^{3}$} is $\epsilon$-EF1 and $\frac{\epsilon}{1+\epsilon}$-MMS. If $e_{3}$ is allocated to agent 2, i.e., $A^{3}_{1}=\{e_{1}\},A^{3}_{2}=\{e_{2},e_{3}\}$, we have $v_{1}(A^{3}_{1})=1=\epsilon\cdot v_{1}(A^{3}_{2}\setminus\{e\})$ for any $e\in A^{3}_{2}$. Thus, $\bf A^{3}$} is $\epsilon$-EF1 and $\epsilon$-MMS.
Therefore, in the above instance, no algorithm can guarantee an NW allocation that is $\alpha$-EF1 or $\alpha$-MMS. This result can be generalized to more than two agents, and we defer the proof to Proposition 1. ∎
Proposition 1.
For general additive functions, no deterministic online algorithm can compute an NW allocation that guarantees $\alpha$-EF1 or $\alpha$-MMS for any $\alpha>0$ when there are more than two agents with additive personalized tri-valued valuation functions.
Proof.
$e_{1}$ $e_{2}$ $\ldots$ $e_{n-2}$ $e_{n-1}$ $e_{n}$ $e_{n+1}$
agent 11 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\ldots$ $\frac{1}{\epsilon}$ 11 $\epsilon$ $\frac{1}{\epsilon}$
agent 22 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\ldots$ $\frac{1}{\epsilon}$ 11 $\epsilon$ $\frac{1}{\epsilon}$
$\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$
agent n2n-2 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\ldots$ $\frac{1}{\epsilon}$ 11 $\epsilon$ $\frac{1}{\epsilon}$
agent n1n-1 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\ldots$ $\frac{1}{\epsilon}$ 11 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
agent $n$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\ldots$ $\frac{1}{\epsilon}$ 11 $\epsilon$ $\frac{1}{\epsilon}$
Table 5: The impossibility results of EF1 and MMS for additive personalized tri-valued valuations when there are more than two agents.
We further extend the impossibility results of approximate EF1 and MMS for $n\geq 3$. Consider the following instance with $n\geq 3$ agents, where the value of each item is shown in Table 5. In the first (n2)(n-2) rounds, to avoid 0-EF1 or 0-MMS, each item is allocated to the agent with the empty bundle. For the item that arrives before round t=n2t=n-2, it has the value of $\frac{1}{\epsilon}$ for any agent. Without loss of generality, we assume that agent $i$ picks the item in the round $t=i$, where $i\leq n$. In round t=n1t=n-1, the arriving item has the value of 1 for any agent, and we assume that it is allocated to agent n1n-1. In round $t=n$, the arriving item has the value of $\epsilon$ for agent $i\in N\setminus\{n-1\}$ and $\frac{1}{\epsilon}$ for agent n1n-1, and it should be allocated to agent $n$ to avoid 0-EF1 or 0-MMS. Finally, in the round $t=n+1$, the item with the value of $\frac{1}{\epsilon}$ for each agent arrives. Due to non-wastefulness, this item must be allocated to one agent. If this item is allocated to agent $i\in N\setminus\{n-1,n\}$, it is not hard to see that the allocation is $\epsilon^{2}$-EF1 and $\frac{\epsilon}{1+\epsilon}$-MMS. If this item is allocated to agent n1n-1, this allocation is $\epsilon$-EF1 and $\frac{\epsilon}{1+\epsilon}$-MMS. If this item is allocated to agent $n$, we can derive that the competitive ratio of EF1 is $\epsilon$ and MMS is $\epsilon$. Combining the above cases, we can conclude that the upper bound of EF1 is $\epsilon$ and MMS is $\epsilon$. For given $\alpha>0$, let $\epsilon<\alpha$, we can conclude that no online algorithm can guarantee an NW allocation that is $\alpha$-EF1 or $\alpha$-MMS for any given $\alpha>0$. ∎
By Theorem 1, we cannot make further progress in the approximations of EF1 and MMS for additive personalized tri-valued valuation functions and beyond in the online setting. In the following parts, we focus on two classes of valuation functions that are more general than additive binary functions, including submodular binary and additive personalized bi-valued valuation functions, and study whether there are positive results about the approximations of EF1 and MMS.
@@ 3.2 Submodular Binary Valuations
This part focuses on submodular binary valuation functions, where for each agent $i\in N$, $v_{i}$ is submodular and the marginal value of any item is 0 or 1. For a subclass of submodular binary valuation functions based on maximum (unweighted) bipartite matching, also known as assignment Munkres (1957) or OXS Lehmann et al. (2001) valuations, Hosseini et al. (2024) show that online allocations satisfying $\frac{1}{2}$-EF1 and $\frac{1}{2}$-MMS exist.
Limitation of Hosseini et al. (2024): Their $\frac{1}{2}$-EF1/MMS guarantees are restricted to OXS valuations, where items must be assigned within predefined groups under rigid matching constraints. Each member in the group can only receive at most one item, and once someone gets an item, it cannot be changed. Thus, their algorithm can only handle the OXS valuation function and its subclasses.
Moving beyond the online allocation model based on matching, we consider the general online allocation problem. Building upon the effective greedy technique in the submodular optimization and the sequential allocating technique to guarantee fairness in Hosseini et al. (2024), we propose the Marginal-Greedy Algorithm (Algorithm 2), greedily assigning each arriving item to the agent who receives the maximum marginal value, for submodular binary valuation functions, and then show that the Marginal-Greedy Algorithm computes an NW allocation that is $\frac{1}{2}$-EF1, $\frac{1}{2}$-MMS, and $\frac{1}{2}$-max-USW.
Output: An NW approximate EF1 and approximate MMS allocation $\bf A$}
1 Initialize: let $\pi=(\pi_{1},\ldots,\pi_{n})$ be an arbitrary order of $n$ agents;
2Let ${\bf A}=(\emptyset,\ldots,\emptyset)$;